If Jamar can run 3/5 of a mile in 2 minutes 30 seconds, what is his rate in miles per minute?
Please help!
Thanks

Answer by gonzo(654)

Rate times Time = Distance
equation is R * T = D
D = 3/5 of a mile.
T = 2 minutes 30 seconds.
R = ?
substituting in the equation.
3/5 miles = R * (2 minutes 30 seconds).
2 minutes 30 seconds = 2.5 minutes.
3/5 miles = R * (2.5 minutes).
R = (3/5 miles) / (2.5 minutes).
3/5 = .6
R = .6 miles / 2.5 minutes.
dividing by 2.5
R = .24 miles / 1 minute = .24 miles per minute.
-----
answer is .24 miles per minute.